OpenStudy (anonymous):

Can somene teach me about how to do square roots like changing it to a radicand and a number? Im really confuse... can someone give me a lesson on it?

OpenStudy (lgbasallote):

can you give an example problem?

OpenStudy (anonymous):

square root of 45....can u explain how to do it?

OpenStudy (lgbasallote):

i see... my method is to express 45 as products of prime numbers.. do you understand what that meanS?

OpenStudy (anonymous):

yes i know what that means. do you use something called a prime factorization tree? or explain your way of doing it(:

OpenStudy (lgbasallote):

well the tree can be useful...is that your teacher's method?

OpenStudy (anonymous):

yeah. but if you have a different way that is easier... could you show me howto do it?

OpenStudy (lgbasallote):

well then i think it'll be the same \[45 = 3 \times 3 \times 5\] \[\sqrt{45} = \sqrt{(3)(3)(5)}\] now..since 3 occurred twice, we can pull it out of the square root giving us \(3\sqrt 5\)

OpenStudy (anonymous):

so when it occurs twice we pull it out? what if there are two numbers that occure twice? maybe you could think of another example to show that(:

OpenStudy (lgbasallote):

\[\sqrt{48}\] \[48 = 2 \times 2 \times 2 \times \times 2 \times 3\] \[\sqrt{48} = \sqrt{(2)(2)(2)(2)(3)}\] 2 occurred twice so pull it out \[2\sqrt{\cancel{(2)(2)}(2)(2)(3)}\] another 2 occurred twice soo \[2(2)\sqrt{\cancel{(2)(2)(2)(2)}(3)} = 4\sqrt 3\]

OpenStudy (anonymous):

so if the number occurs more than twice then we multiply them together like how you multiplied 2 and 2 to get 4?

OpenStudy (lgbasallote):

yep

OpenStudy (lgbasallote):

can you find the \(\sqrt{24}\)?

OpenStudy (anonymous):

\[4\sqrt{3}\] i think?

OpenStudy (lgbasallote):

but that's the answer to sqrt 48 :P i knew you'd get confused \[\sqrt{24} = \sqrt{(2)(2)(2)(3)}\] 2 occurred twice so pull it out \[2\sqrt{\cancel{(2)(2)}(2)(3)}\] note that no other number appeared TWICE so \(2\sqrt 6\) is the final answer

OpenStudy (anonymous):

oh ok. thanks for the clarification. give me one more then ill b good i think

OpenStudy (anonymous):

nevermind i g2g. i got it